Buffalo BBQ Restaurant is trying to become more efficient in training its chefs. It is experimenting with two training programs aimed at this objective. Both programs have basic and advanced training modules. The restaurant has provided the following data regarding the two programs after two weeks of implementation:

Training Program A Training Program B
New chef # 1 2 3 4 5 6 7 8 9 10
Hours of basic training 20 24 26 22 25 24 25 30 30 24
Hours of advanced training 10 6 9 11 11 3 2 0 2 3
Number of chef mistakes 10 14 17 12 12 9 6 8 5 7

a. Compute the following performance metrics for each program:

(1) Average hours of employee training per chef, rounded to one decimal place.

Program A:  hrs. per chef

Program B:  hrs. per chef

(2) Average number of mistakes per chef, rounded to one decimal place.

Program A:  mistakes per chef

Program B:  mistakes per chef

b. Which program should the restaurant implement moving forward?
Answer #1

Average hours of employee training per chef,

  1. Program A, 32.8 hours per chef.
    1. Total training hours = 164,
    2. Number of chefs = 5,
    3. Average hours = 164 ÷5 = 32.8.
  2. Program B, 28.6 hours per chef.
    1. Total trainin hours = 143,
    2. Number of chefs = 5,
    3. Average hours = 143÷5 = 28.6.

Average number of mistakes per chef,

  1. Program A, 13 mistakes per chef.
    1. Total mistakes = 65,
    2. Number of chefs = 5,
    3. Average = 65 ÷ 5 = 13.
  2. Program B, 7 mistakes per chef.
    1. Total mistakes = 35,
    2. Number of chefs = 5,
    3. Average = 35÷5 = 7.

The restaurant should implement program B because, the number of training hours are less and even though they are less, the number of mistakes committed by chefs are less in program B. Therefore the overall costs will reduce.